find the equation of the line passing through point (-1,-2) and is parallel to the line joining the points a(3,4) and b(7,8).
 

Grade:11

3 Answers

jawwad
8 years ago
x-y-1=0
sneha
8 years ago
i need the steps please
 
sneha
8 years ago
m=y2-y1/x2-x1 
m=1=tan 45’
m=8-4/7-3=1
y-y1=m(x-x1)
y+2=(1)(x+1)
so, the equation is x-y-1=0